OpenStudy (safia21):

Identify the parent function for g from its function rule. The graph g on your calulator and describe what transformation of the parent function represents gx)= x^2- 1 gx)= square root of x-2 gx) = x^3 + 3

OpenStudy (anonymous):

g(x)=x^2 g(x)=squareroot of x g(x)=x^3

OpenStudy (anonymous):

those are the parent functions

OpenStudy (anonymous):

duh

OpenStudy (safia21):

haha thanks?

OpenStudy (anonymous):

yw